Solve the equation.
2/3y = 4
Any help with this equation would be Greatly appreciated.

Answer by checkley71(8403)

2Y/3=4 NOW CROSS MULTIPLY
2Y=3*4
Y=12/2
Y=6 ANSWER.
PROOF
2*6/3=4
12/3=4
4=4